# Scalable Image Coding for Humans and Machines Using Feature Fusion Network

### v3

@@ I Introduction
Figure 1: Image processing flow of the Scalable Image Coding method. (a): conventional method assuming a specific image recognition model, (b): proposed method assuming various image recognition models.
@@ II-B Channel-Conditional Learned Image Compression
Figure 2: Examples of the mask image. (a) : original image in COCO dataset, (b) : mask image generated using Segment Anything.
Figure 3: Image processing flow of the proposed scalable image coding method. The LIC model in the upper row is the SA-ICM, which is used to compress images for machines. The LIC model in the lower is an additional information compression model, which converts an coded image for machine into an image for human vision.
Figure 4: Model structure of the proposed feature fusion network. The purple and pink squares represent the features of the image compression model for machines and that of the additional information compression model, respectively.
@@ II-C SA-ICM
SA-ICM is an LIC model that learns only the edge information of images, with a model structure is identical to the LIC model proposed by J. Liu et al.[35]. By learning the edge information produced with Segment Anything [36], we can create an LIC model capable of decoding only the edges of an image. Meanwhile, the decoded image retains the color of the original image as well as the edge information of the object and background, making the image suitable for image recognition models. This model has been proven to be an effective image compression method for object detection models[37], object segmentation models[38], and background segmentation models[39]. The usefulness of SA-ICM is demonstrated using three different datasets[40, 41, 42]. The loss function for training this model is expressed by the following equation:
@@ III Proposed Method
Figure 5: Example of coded image. (a) : original image. (b) : decoded images for machines using SA-ICM. (c) : decoded images for human vision using additional information.
@@ IV-A Performance in Image Compression for Humans
Figure 7: Relationship between the parameter of $m$ and compression performance in the proposed method.
@@ IV-B Parameter Reduction in the Additional Information Compression Model
As noted in Section 3.2, we investigate the impact of parameter reduction in the additional information compression model. In this experiment, the parameter $n$ is fixed at 5, consequently implying that the value of $m$ falls within the range of greater than 1 and less than 5. The size of the model can be adjusted based on the value of $m$. Setting a larger value for $m$ increases the number of intermediate features. Hence, the number of parameters also increases. The relationship between the value of $m$ and the number of parameters of the additional information compression model is shown in Table I. The relationship between the value of $m$ and the image quality of the decoded image is shown in Table II and Fig. 7. From these results, it can be seen that good quality images can be decoded without much dependence on the value of $m$. Therefore, depending on the value of $m$, it is possible to reduce the number of parameters while maintaining the quality of the decoded image.

### v4

@@ I Introduction
Figure 1: Image processing flow of the scalable image coding method. (a): Conventional method aimed for a specific image recognition model. (b): Proposed method compatible for various image recognition models.
@@ II-B Channel-Conditional Learned Image Compression
Figure 2: Examples of the mask image. (a) : Original image in COCO dataset. (b) : Mask image generated using Segment Anything Model.
Figure 3: The processing flow of the proposed scalable image coding method. The LIC model in the upper row is SA-ICM, which compresses images for machines. The LIC model in the lower is an additional information compression model, which converts coded images for machines into images for humans.
Figure 4: Model structure of the feature fusion network. The purple and pink squares represent the features of the image compression model for machines and that of the additional information compression model, respectively.
@@ II-C SA-ICM
SA-ICM is an LIC model that learns only the edge information of images, with a model structure is identical to the LIC model proposed by J. Liu et al.[35]. By learning the edge information produced with Segment Anything [36], we can create an LIC model capable of decoding only the edges of an image. Meanwhile, the decoded image retains the color of the original image as well as the edge information of the object and background, making the image suitable for image recognition models. This model has been proven to be an effective image compression method for object detection models[37], object segmentation models[38], and background segmentation models[39]. The usefulness of SA-ICM is demonstrated using three different datasets[40, 41, 42]. The loss function is expressed by the following equation:
@@ III Proposed Method
Figure 5: Example of compressed images for humans and machines. (a) : original image. (b) : decoded images for machines using SA-ICM. (c) : decoded images for human vision using additional information.
@@ IV-A Performance in Image Compression for Humans
Figure 7: Relationship between the change in parameter $m$ and compression performance of our proposed method.
@@ IV-B Parameter Reduction in the Additional Information Compression Model
As noted in Section 3.2, we investigate the impact of parameter reduction in the additional information compression model. In this experiment, the parameter $n$ is fixed at 5, consequently implying that the value of $m$ falls within the range of greater than 1 and less than 5. The size of the model can be adjusted based on the value of $m$. Setting a larger value for $m$ increases the number of intermediate features. Hence, the number of parameters also increases. The relationship between the value of $m$ and the number of parameters of the additional information compression model is shown in Table I. The relationship between the value of $m$ and the image quality of the decoded image is shown in Table II and Fig. 7. These results indicate that decoding images for humans does not heavily rely on the value of $m$, particularly at low bitrates. For instance, at bitrates around 0.6 [bpp], images of nearly identical quality are decoded whether $m$ is 3, 4, or 5. Therefore, depending on the value of $m$, it is possible to reduce the number of parameters while maintaining the quality of the decoded image.